# ReST-MCTS: LLM Self-Training via Process Reward Guided Tree Search

### v2

@@ Abstract
11footnotetext: DZ and SZ contributed equally. 22footnotetext: Work done while DZ visited at Caltech.33footnotetext: Work done while SZ interned at Zhipu AI.
@@ 1 Introduction
Large Language Models (LLMs) are mostly trained on human-generated data. But as we approach the point where most available high-quality human-produced text on the web has been crawled and used for LLM training [1], the research focus has shifted towards using LLM-generated content to conduct self-training [2, 3, 4, 5, 6, 7]. Similar to most Reinforcement Learning (RL) problems, LLM self-training requires a reward signal. Most existing reinforced self-improvement approaches (e.g., STaR [4], RFT [5], $\text{ReST}^{\text{EM}}$ [6], V-STaR [7]) assume to have access to a ground-truth reward model (labels from supervised dataset, or a pre-trained reward model). These approaches use an LLM to generate multiple samples for each question, and assume the one that leads to high reward (correct solution) is the high-quality sample, and later train on these samples (hence self-training). Such procedures can be effective in improving LLM performance, in some cases solving reasoning tasks that the base LLM cannot otherwise solve [8, 9, 10].
@@
However, a key limitation of the above procedure is that even if a reasoning trace results in a correct solution, it does not necessarily imply that the entire trace is accurate. LLMs often generate wrong or useless intermediate reasoning steps, while still finding the correct solution by chance [17]. Consequently, a self-training dataset can often contain many false positives — intermediate reasoning traces or plans are incorrect, but the final output is correct — which limits the final performance of LLM fine-tuning for complex reasoning tasks [18, 19]. One way to tackle this issue is to use a value function or reward model to verify reasoning traces for correctness (which then serves as a learning signal for self-training) [1, 12]. However, training a reliable reward model to verify every step in a reasoning trace generally depends on dense human-generated annotations (per reasoning step) [1], which does not scale well. Our research aims to address this gap by developing a novel approach that automates the acquisition of reliable reasoning traces while effectively utilizing reward signals for verification purposes. Our key research question is: How can we automatically acquire high-quality reasoning traces and effectively process reward signals for verification and LLM self-training?
@@ 2 Background on Reasoning & Self-Training
Tree-Search & Value Function. Another idea is to use tree-structured reasoning traces [24, 14], that branch from intermediate reasoning steps. One key issue in using a so-called tree-search reasoning algorithm is the need to have a value function to guide the otherwise combinatorially large search process [14]. Two common value functions include Outcome Reward Models (ORMs) [25], which are trained only on the correctness of the final answer, and Process Reward Models (PRMs) [1], which are trained on the correctness of each reasoning step. We assume $r_{s_{k}}$ is the PRM’s output sigmoid score at $k$-th step. Our ReST-MCTS approach uses tree-search to automatically learn a good PRM.
@@
Self-Training. At a high level, there are two steps to self-training [6, 12]. The first step is generation, where we sample multiple reasoning traces using $\pi$ (in our case, tree-structured traces). The second step is improvement, where a learning signal is constructed on the reasoning traces, which is then used to fine-tune $\pi$. The process can repeat multiple iterations.
@@ 3.1 Search-based Reasoning Policy for LLM
Quality Value $v_{k}$ for a Partial Solution. The value or process reward $v_{k}$ of the partial solution $p_{k}=[s_{1},s_{2},\cdots,s_{k}]$ should satisfy the following basic qualities:
@@
  • Reflecting probability of correctness: $v_{k}$ serves as a reflection of the probability that a partial solution is a complete and correct answer. Higher values of $v_{k}$ indicate better quality or a higher likelihood of being closer to a correct answer.
@@
The weighted reward $w_{s_{k}}$ of the current step provides a measure of the quality and contribution of that specific step towards the overall solution. Based on $m_{k}$ (where $m_{k}=K-k$ and $K$ is the total number of reasoning steps of a solution $s$), previous quality value $v_{k-1}$, and $r_{s_{k}}$, we can update the definition of the weighted reward $w_{s_{k}}$ iteratively as follows:
@@
3.2 Self-Training ReST-MCTS Pipeline
@@ 4 Experiments
$\bullet$ Process Reward models which are compared with the state-of-the-art techniques, such as MATH-SHEPHERD (MS) and SC + MS on GSM8K and MATH500, as shown in Table 3. Results indicate ReST-MCTS learns a good PRM and our reward model implements higher accuracy.
@@ 4.1 Initialization of Value Model
To obtain accurate feedback from the environment, we build the value model’s initial train set $D_{V_{0}}$ from a set of selected science or math questions $D_{0}$ using process reward (value) inference, with no human labeling process required. Then, we finetune the ChatGLM3-6B [27, 28] and Mistral-7B [29] model on this dataset, respectively, obtaining initial value models that, as variants of PRM, guide the LLM tree search for higher-quality solutions upon both math and science questions.
Fine-grained dataset for science and math. Aiming to gather value train data for science, we integrate questions of a lean science dataset $D_{sci}$ within SciInstruct [10] into $D_{0}$. This dataset consists of 11,554 questions, where each question is paired with a correct step-by-step solution. For each question $q^{(i)}(i=1,2,\cdots,N)$ and corresponding solution $s^{(i)}=s^{(i)}_{1,2,\cdots,K_{i}}$ in $D_{sci}$, we extract all partial solutions to form samples $d_{k}^{(i)}=[q^{(i)},s^{(i)}_{1,2,\cdots,k}(p_{k}^{(i)})](k=1,2,\cdots,K_{i})$. To make the value model distinguish false steps, we also employ a LLM policy (ChatGLM2) that is basically incompetent for reasoning tasks of this difficulty to generate single steps $s_{k+1}^{(i)^{\prime}}$ given $q^{(i)}$ and $p_{k}^{(i)}$, obtaining new partial solutions $p_{k+1}^{(i)^{\prime}}=[s_{1,2,\cdots,k}^{(i)},s_{k+1}^{(i)^{\prime}}]$ and new samples $d^{(i)}_{k,j}=[q^{(i)},s^{(i)}_{1,2,\cdots,k},s_{k+1,j}^{(i)^{\prime}}](j=1,2,3)$. For simplicity, the generated steps are regarded as incorrect. Afterward, we derive target quality values for all samples $d^{(i)}_{k,j}$ and $d_{k}^{(i)}$ and use them to construct $D_{V_{0}}$, which is illustrated in Appendix B.1. We adopt an alternative method to generate value train data for math, as shown in Appendix B.1.
@@ 4.2 Evaluating Self-Improvement of ReST-MCTS
$\bullet$ Iterative performance improvement on reward model. We also compare how our iterative trained policy and value model can improve the overall search results under the same token usage on the test set of MATH [33]. See implementation details in Appendix E.3. We show results in Figure 2 (a), where ReST-MCTS (Iter #1) greatly outperforms most baselines but does not completely surpass Self-Consistency. In comparison, after more iterations of self-train, verification based on the enhanced value model basically outperforms Self-Consistency on every point, achieving the highest accuracy of $48.5\%$ that significantly exceeds the $42.5\%$ of Self-Consistency. This indicates the effectiveness of our self-training pipeline.
@@ 4.3 Evaluating Reward Guidance and Reasoning Policy of ReST-MCTS
Performance comparison of various verification models. As [1] suggested, different value models or reward models vary in accuracy and fineness. We perform tests on the questions of the GSM8K and MATH500 using multiple rewards (value) models and verification methods. It is worth noting that we include the same experiment settings of MATH-SHEPHERD (MS) [12] as a comparison since it also adopts an automatic train data generation method for reward models. For SC+ReST-MCTS, we utilize the same CoT-based sampling strategy as MS, except that SC is performed according to our own value model’s output rather than the reward model of MS, which makes this a direct comparison of different reward model training approaches. We record the model accuracy of Mistral-7B: MetaMATH on the selected test set, which is as shown in Table 3. Results indicate that compared to MS and SC+MS, SC+ReST-MCTS (Value) exhibits higher improvement in solution accuracy on both GSM8K and MATH. This confirms the effectiveness of our value model, further indicating that our definitions of quality value and weighted reward are valid or possibly even better.
Performance comparison under the same search budget. Though the MCTS-based search methods demonstrate significant improvement in model performance, they often require a considerable amount of token input and completion, which makes it quite costly in some circumstances. Therefore, we conduct more experiments to investigate the relationship between search token budget and model performance on science questions selected from SciBench comparing ReST-MCTS and the same baselines employed for MATH, which are elaborated in Appendix E.3. Since our self-training procedure is primarily conducted on math data, so we do not consider the effects of self-training in this case. However, we point out that this can still be further investigated as a study of transfer learning for self-training paradigms. Figure 2 (b) shows the accuracy of different approaches when the completion budget changes. Results indicate that the ReST-MCTS greatly outperforms other baselines despite insufficient budget. We notice that although CoT-based methods can improve greatly by increasing the sample budget, they tend to quickly converge to a limited accuracy, which is not as satisfying as the ReST-MCTS.
@@ 5.1 Large Language Model Training
Large Language Models (LLMs) [36, 37, 38] have emerged as a notable success in various natural language tasks. Recent studies focus on improving the reasoning capabilities of LLMs, including collecting high-quality or larger domain-specific data [39, 40, 41, 42, 10, 43], designing elaborate prompting [22, 44, 45, 46], or training supervised learning [10, 31, 32, 47] or reinforcement learning (RL) [48, 49, 50, 16]. When LLMs are trained with the RL algorithm, the generation of LLMs can be naturally expressed as the Markov Decision Process (MDP) and optimized for specific objectives. According to this formula, InstructGPT [51] has achieved remarkable success in optimizing LLMs to align human preferences by utilizing RL from Human Feedback (RLHF) [52]. RLAIF then uses AI feedback to extend RL from human feedback [53].
@@ 5.2 Large Language Model Reasoning
LLM reasoning algorithms include prompt-based chain-of-thought (CoT) [22], planning-based represented by tree-of-thought (ToT) [24]. Scientific reasoning has several categories to mine the potential of existing large language models, resulting from different performances for problem-solving. Previous studies have attempted to outperform the direct generation. For example, in this paper [54], an approach for generating solutions in a step-by-step manner is proposed, another model or function is used to select the top-ranked answers, and hallucination is avoided by limiting the output to a narrower set. [55] presents a maieutic prompting inference method, which can generate abductive explanations of various hypotheses explained by recursion, eliminate contradicting candidates, and achieve logically consistent reasoning. Chain-of-thoughts (CoT) [22] imitates the thought process like humans to provide step-by-step solutions given a question. Self-Consistency CoT [23] improves the reliability and Self-Consistency of answers by sampling multiple interpretations from LM and then selecting the final answer that appears most frequently. Tree-of-Thoughts (ToT) [24] further generalizes the CoT methodology by considering multiple different reasoning paths in the tree and exploring coherent units of thought to execute thoughtful decision-making. In our work, we benchmark hard science reasoning tasks [34, 35] against [22, 24].
@@ C.2 Data Generation Process and Specific Example for Reward Inference
The data generation process of our self-training approach consists of mainly $4$ stages, namely search, prune, verify, and reward inference, which is demonstrated in Figure 4. For reward inference, a detailed example is shown in Figure 3.
@@ E.1 Training and Evaluation of Initial Value Model
$\vspace{-0.1cm}Accuracy=\frac{1}{t}\Sigma_{i=1}^{t}\ I(|clip(V_{\theta}(p_{i}|q_{i}),0,1)-v_{i}^{*}|<0.1)$ (40)
@@ E.2 Benchmark Setup
To compare the performance of different search methods, we construct a standardized benchmark test that can be generally used on labeled science or math datasets like MATH, SciBench, and SciEval. Aside from the ReST-MCTS, we incorporate two other baselines: chain-of-thought (CoT) and tree-of-thought (ToT). For each method, specialized prompts $P$ are designed to execute the search process. Besides, an inference model $\pi$ and value model $V$ are deployed to provide deduction and feedback. Concerning the CoT baseline, we use Self-Consistency to calculate accuracy. For the ToT baseline, we use a simple greedy depth-first search (DFS) algorithm with node values assigned by the value model. The algorithm stops exploitation when a max depth of $10$ is reached and ends when a node value exceeds the threshold $0.9$. For ReST-MCTS, self-critic is used and the ending threshold is also set to $0.9$. The rollout step limit $m$ is set to $2$, $\alpha$ is set to $0.5$, and the number of iterations $T$ is set to $50$ by default. Moreover, both tree search algorithms use $b=3$ by default, where $b$ is the number of samples generated in the expansion process as mentioned in the former sections. After the search process, the policy is prompted to extract the final answer based on the obtained solution, which is then compared with the ground truth to determine correctness. The results of these methods on benchmarks are illustrated in Section 4.3.
@@ E.3 Baselines of Search Verification
  • PRM + Best-of-N (BoN) With value model $V_{\theta}$ used as PRM, we perform DFS-based tree search. Every selected solution $s$ is evaluated by a PRM score $r_{PRM}=\Pi_{i=1}^{K}v_{i}$. The one with the highest PRM score among all $N$ solutions is regarded as the final output. Under this setting, $b$ is set to $3$, while $N$ is used to control token usage.

### v3

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Work done while DZ visited at Caltech.
@@ 1 Introduction
[…]
@@
However, a key limitation of the above procedure is that even if a reasoning trace results in a correct solution, it does not necessarily imply that the entire trace is accurate. LLMs often generate wrong or useless intermediate reasoning steps, while still finding the correct solution by chance [17]. Consequently, a self-training dataset can often contain many false positives — intermediate reasoning traces or plans are incorrect, but the final output is correct — which limits the final performance of LLM fine-tuning for complex reasoning tasks [18; 19]. One way to tackle this issue is to use a value function or reward model to verify reasoning traces for correctness (which then serves as a learning signal for self-training) [1; 12]. However, training a reliable reward model to verify every step in a reasoning trace generally depends on dense human-generated annotations (per reasoning step) [1], which does not scale well. Our research aims to address this gap by developing a novel approach that automates the acquisition of reliable reasoning traces while effectively utilizing reward signals for verification purposes. Our key research question is: How can we automatically acquire high-quality reasoning traces and effectively process reward signals for verification and LLM self-training?
@@ 2 Background on Reasoning & Self-Training
Tree-Search & Value Function. Another idea is to use tree-structured reasoning traces [24; 14], that branch from intermediate reasoning steps. One key issue in using a so-called tree-search reasoning algorithm is the need to have a value function to guide the otherwise combinatorially large search process [14]. Two common value functions include Outcome Reward Models (ORMs) [25], which are trained only on the correctness of the final answer, and Process Reward Models (PRMs) [1], which are trained on the correctness of each reasoning step. We assume $r_{s_{k}}$ is the PRM’s output sigmoid score at $k$-th step. Our ReST-MCTS approach uses tree-search to automatically learn a good PRM.
@@
Self-Training. At a high level, there are two steps to self-training [6; 12]. The first step is generation, where we sample multiple reasoning traces using $\pi$ (in our case, tree-structured traces). The second step is improvement, where a learning signal is constructed on the reasoning traces, which is then used to fine-tune $\pi$. The process can repeat for multiple iterations.
@@ 3.1 Search-based Reasoning Policy for LLM
Value $v_{k}$ for a Partial Solution. The value (process) reward $v_{k}$ of the partial solution $p_{k}=[s_{1},s_{2},\cdots,s_{k}]$ should satisfy the following basic qualities:
@@
  • Reflecting probability of correctness: $v_{k}$ reflects the probability that a partial solution is a complete and correct answer. Higher values of $v_{k}$ indicate better quality or a higher likelihood of being closer to a correct answer.
@@
The weighted reward $w_{s_{k}}$ of the current step provides a measure of the quality and contribution of that specific step towards the overall solution. Based on $m_{k}$ (where $m_{k}=K-k$ and $K$ is the total number of reasoning steps of a solution $s$), previous quality value $v_{k-1}$, and $r_{s_{k}}$ in MATH-SHEPHERD [12], we can update the definition of the weighted reward $w_{s_{k}}$ iteratively as follows:
@@
3.2 Self-Training Pipeline
@@ 4 Experiments
$\bullet$ Process reward models which are compared with the state-of-the-art techniques, such as MATH-SHEPHERD (MS) and SC + MS on GSM8K and MATH500, as shown in Table 3. Results indicate that the ReST-MCTS learns a good PRM and our reward model implements higher accuracy.
@@ 4.1 Initialization of Value Model
To obtain accurate feedback from the environment, we build the value model’s initial train set $D_{V_{0}}$ from a set of selected science or math questions $D_{0}$ using process reward (value) inference, with no human labeling process required. Then, we finetune the ChatGLM3-6B [27; 28] and Mistral-7B [29] model on this dataset, respectively, obtaining initial value models that, as variants of PRM, guide the LLM tree search for higher-quality solutions upon both math and science questions.
Fine-grained dataset for science and math. Aiming to gather value train data for science, we integrate questions of a lean science dataset $D_{sci}$ within SciInstruct [10] into $D_{0}$. This dataset consists of 11,554 questions, where each question is paired with a correct step-by-step solution. For each question $q^{(i)}(i=1,2,\cdots,N)$ and corresponding solution $s^{(i)}=s^{(i)}_{1,2,\cdots,K_{i}}$ in $D_{sci}$, we extract all partial solutions to form samples $d_{k}^{(i)}=[q^{(i)},s^{(i)}_{1,2,\cdots,k}(p_{k}^{(i)})](k=1,2,\cdots,K_{i})$. To make the value model distinguish false steps, we also employ a LLM policy (ChatGLM2) that is basically incompetent for reasoning tasks of this difficulty to generate single steps $s_{k+1}^{(i)^{\prime}}$ given $q^{(i)}$ and $p_{k}^{(i)}$, obtaining new partial solutions $p_{k+1}^{(i)^{\prime}}=[s_{1,2,\cdots,k}^{(i)},s_{k+1}^{(i)^{\prime}}]$ and new samples $d^{(i)}_{k,j}=[q^{(i)},s^{(i)}_{1,2,\cdots,k},s_{k+1,j}^{(i)^{\prime}}](j=1,2,3)$. For simplicity, the generated steps are regarded as incorrect. In total, we collect $473.4$k samples for training the initial value model. Afterward, we derive target quality values for all samples $d^{(i)}_{k,j}$ and $d_{k}^{(i)}$ and use them to construct $D_{V_{0}}$, which is illustrated in Appendix B.1. We adopt an alternative method to generate value train data for math, as shown in Appendix B.1.
@@ 4.2 Evaluating Self-Improvement of ReST-MCTS
$\bullet$ Iterative performance improvement on reward model. We also compare how our iterative trained policy and value model can improve the overall search results under the same token usage on the test set of MATH [33]. See implementation details in Appendix E.3. We show results in Figure 2 (a), where ReST-MCTS (Iter #1) greatly outperforms most baselines but does not completely surpass Self-Consistency. In comparison, after more iterations of self-training, verification based on the enhanced value model basically outperforms Self-Consistency on every point, achieving the highest accuracy of $48.5\%$ that significantly exceeds the $42.5\%$ of Self-Consistency. This indicates the effectiveness of our self-training pipeline.
@@ 4.3 Evaluating Reward Guidance and Reasoning Policy of ReST-MCTS
Performance Comparison of Various Verification Models. As [1] suggested, different value models or reward models vary in accuracy and fineness. We perform tests on the questions of the GSM8K and MATH500 using multiple reward models and verification methods. It is worth noting that we include the same experiment settings of MATH-SHEPHERD (MS) [12] as a comparison since it also adopts an automatic train data generation method for reward models. For SC+ReST-MCTS, we utilize the same CoT-based sampling strategy as MS, except that SC is performed according to our own value model’s output rather than the reward model of MS, which makes this a direct comparison of different reward model training approaches. We record the model accuracy of Mistral-7B: MetaMATH on the selected test set, which is as shown in Table 3. Results indicate that compared to MS and SC+MS, SC+ReST-MCTS (Value) exhibits higher improvement in solution accuracy on both GSM8K and MATH. This confirms the effectiveness of our value model, further indicating that our definitions of quality value and weighted reward are valid or possibly even better.
Performance Comparison under the Same Search Budget. Though the MCTS-based search methods demonstrate significant improvement in model performance, they often require a considerable amount of token input and completion, which makes it quite costly in some circumstances. Therefore, we conduct more experiments to investigate the relationship between search token budget and model performance on science questions selected from SciBench comparing ReST-MCTS and the same baselines employed for MATH, which are elaborated in Appendix E.3. Since our self-training procedure is primarily conducted on math data, so we do not consider the effects of self-training in this case. However, we point out that this can still be further investigated as a study of transfer learning for self-training paradigms. Figure 2 (b) shows the accuracy of different approaches on SciBench when the completion budget changes. Results indicate that the ReST-MCTS greatly outperforms other baselines despite insufficient budget. We notice that although CoT-based methods can improve greatly by increasing the sample budget, they tend to quickly converge to a limited accuracy, which is not as satisfying as the ReST-MCTS.
@@ 5.1 Large Language Model Training
Large Language Models (LLMs) [36; 37; 38] have emerged as a notable success in various natural language tasks. Recent studies focus on improving the reasoning capabilities of LLMs, including collecting high-quality or larger domain-specific data [39; 40; 41; 42; 10; 43], designing elaborate prompting [22; 44; 45; 46], or training supervised learning [10; 31; 32; 47] or reinforcement learning (RL) [48; 49; 50; 16]. When LLMs are trained with the RL algorithm, the generation of LLMs can be naturally expressed as the Markov Decision Process (MDP) and optimized for specific objectives. According to this formula, InstructGPT [51] has achieved remarkable success in optimizing LLMs to align human preferences by utilizing RL from Human Feedback (RLHF) [52]. RLAIF then uses AI feedback to extend RL from human feedback [53]. Our work aims to propose an LLM self-training method via process rewards guided tree search.
@@ 5.2 Large Language Model Reasoning
LLM reasoning algorithms include prompt-based chain-of-thought (CoT) [22], planning-based represented by tree-of-thought (ToT) [24]. Scientific reasoning has several categories to mine the potential of existing large language models, resulting from different performances for problem-solving. Previous studies have attempted to outperform the direct generation. For example, in this paper [54], an approach for generating solutions in a step-by-step manner is proposed, another model or function is used to select the top-ranked answers, and hallucination is avoided by limiting the output to a narrower set. [55] presents a maieutic prompting inference method, which can generate abductive explanations of various hypotheses explained by recursion, eliminate contradicting candidates, and achieve logically consistent reasoning. Chain-of-thoughts (CoT) [22] imitates the thought process like humans to provide step-by-step solutions given a question. Self-Consistency CoT [23] improves the reliability and Self-Consistency of answers by sampling multiple interpretations from LM and then selecting the final answer that appears most frequently. Tree-of-Thoughts (ToT) [24] further generalizes the CoT methodology by considering multiple different reasoning paths in the tree and exploring coherent units of thought to execute thoughtful decision-making. In our work, we benchmark hard science reasoning tasks against [22; 24; 34; 35].
@@ 6 Conclusion
Acknowledgments
Dan and Sining would like to thank Zhipu AI for sponsoring the computation resources used in this work. Yisong is supported in part by NSF #1918655. Yuxiao and Jie are supported in part by the NSFC 62276148, NSFC for Distinguished Young Scholar 62425601, a research fund from Zhipu, New Cornerstone Science Foundation through the XPLORER PRIZE and Tsinghua University (Department of Computer Science and Technology) - Siemens Ltd., China Joint Research Center for Industrial Intelligence and Internet of Things (JCIIOT). Corresponding authors: Yisong Yue, Yuxiao Dong, and Jie Tang.
Appendix
[…]
@@ C.2 Data Generation Process and Specific Example for Reward Inference
The data generation process of our self-training approach consists of mainly four stages, namely search, prune, verify, and reward inference, which is demonstrated in Figure 4. For reward inference, a detailed example is shown in Figure 3.
@@ E.1 Training and Evaluation of Initial Value Model
$\vspace{-0.1cm}\text{Accuracy}=\frac{1}{t}\Sigma_{i=1}^{t}\ I(|clip(V_{\theta}(p_{i}|q_{i}),0,1)-v_{i}^{*}|<0.1)$ (40)
@@ E.2 Benchmark Setup
To compare the performance of different search methods, we construct a standardized benchmark test that can be generally used on labeled science or math datasets like MATH, SciBench, and SciEval. Aside from the ReST-MCTS, we incorporate two other baselines: chain-of-thought (CoT) and tree-of-thought (ToT). For each method, specialized prompts $P$ are designed to execute the search process. Besides, an inference model $\pi$ and value model $V$ are deployed to provide deduction and feedback. Concerning the CoT baseline, we use Self-Consistency to calculate accuracy. For the ToT baseline, we use a simple greedy depth-first-search (DFS) algorithm with node values assigned by the value model. The algorithm stops exploitation when a max depth of $10$ is reached and ends when a node value exceeds the threshold $0.9$. For ReST-MCTS, self-critic is used and the ending threshold is also set to $0.9$. The rollout step limit $m$ is set to $2$, $\alpha$ is set to $0.5$, and the number of iterations $T$ is set to $50$ by default. Moreover, both tree search algorithms use $b=3$ by default, where $b$ is the number of samples generated in the expansion process as mentioned in the former sections. After the search process, the policy is prompted to extract the final answer based on the obtained solution, which is then compared with the ground truth to determine correctness. The results of these methods on benchmarks are illustrated in Section 4.3.
@@ E.3 Baselines of Search Verification
  • PRM + Best-of-N (BoN) With value model $V_{\theta}$ used as PRM, we perform DFS-based tree search. Every selected solution $s$ is evaluated by a PRM score $r_{\text{PRM}}=\Pi_{i=1}^{K}v_{i}$. The one with the highest PRM score among all $N$ solutions is regarded as the final output. Under this setting, $b$ is set to $3$, while $N$ is used to control token usage.